OpenStudy (mathidiot):

The area of a given hexagon is equal to the area of an equilateral triangle whose perimeter is 36 inches. Find the length of a side of the regular hexagon.

jhonyy9 (jhonyy9):

an equilateral triangle has 3 equal sides so than the perimeter is 36 inches so this mean that a side has a length of 12 inches so than you know that the length of a side of an equilateral triangle is equal 12 so can yo calculi arae of this triangle equal how many ?

jhonyy9 (jhonyy9):

sorry - can you calculi area of this triangle ?

OpenStudy (mathidiot):

base x height?

jhonyy9 (jhonyy9):



jhonyy9 (jhonyy9):

no area of a triangle is base time height divide 2

jhonyy9 (jhonyy9):

but in this case first you need calculi the length of this height - yes ?

OpenStudy (mathidiot):

yes?

jhonyy9 (jhonyy9):

any idea to how you can calculi the length of this height ?

jhonyy9 (jhonyy9):

you can using pythagora"s theorem or the sin of 60 degree - what you like using ?

jhonyy9 (jhonyy9):

or tan or cot ,what you like using

OpenStudy (mathidiot):

pyth

jhonyy9 (jhonyy9):

do you know exactly this i wan writing now

jhonyy9 (jhonyy9):

so what say pyth in this case ?

jhonyy9 (jhonyy9):

height so h^2 = ?

jhonyy9 (jhonyy9):

you dont know it ?

jhonyy9 (jhonyy9):

h^2 = 12^2 -6^2 yes ?

OpenStudy (mathidiot):

yes

jhonyy9 (jhonyy9):

so than h = ?

OpenStudy (mathidiot):

idk

jhonyy9 (jhonyy9):

h = sqrt(144-36) = ?

OpenStudy (mathidiot):

108

jhonyy9 (jhonyy9):

yes but sqrt108 = ?

OpenStudy (mathmate):

If you like to try a different approach, here are some hints: 

OpenStudy (mathmate):

Oops, the bottom line should read S^2/A=s^2/a s^2=S^2(a/A) s^2=S^2(1/6) s=S*sqrt(1/6)=S/sqrt(6)